OpenStudy (anonymous):

Solve by factoring. 16x^2-25=0

OpenStudy (anonymous):

\[16x^{2} = 25\]\[x^{2} = 25/16\]\[x = \sqrt { 25/16 }\]

OpenStudy (anonymous):

\[16x^2-25 \implies 16x^2=25 \implies x^2={25 \over 16} \implies x=\pm{5 \over 4}\]